OpenStudy (anonymous):

The depth d in feet of the water in a bay at is given by d(t) = 3/2 sin (5πt/31) + 23 where t is time. Graph the depth of the water as a function of time. What is the maximum depth of the water to the nearest tenth of a foot?

OpenStudy (anonymous):

A plot of \[\frac{3}{2} \sin \left(\frac{5 \pi t}{31}\right)+23 \]from 0 through 3 pi is attached.